# Persistence of AR($1$) sequences with Rademacher innovations and linear mod $1$ transforms

### v2

@@ Abstract.
We study the probability that an AR(1) Markov chain $X_{n+1}=aX_{n}+\xi_{n+1}$, where $a\in(0,1)$ is a constant, stays non-negative for a long time. We find the exact asymptotics of this probability and the weak limit of $X_{n}$ conditioned to stay non-negative, assuming that the i.i.d. innovations $\xi_{n}$ take only two values ±$\pm 1$ and $a\leq\frac{2}{3}$. This limiting distribution is quasi-stationary. It has no atoms and is singular with respect to the Lebesgue measure when $\frac{1}{2}<a\leq\frac{2}{3}$, except for the case $a=\frac{2}{3}$ and $\mathbb{P}(\xi_{n}=1)=\frac{1}{2}$, where this distribution is uniform on the interval $[0,3]$. This is similar to the properties of Bernoulli convolutions.
It turns out that for the ±$\pm 1$ innovations there is a close connection between $X_{n}$ killed at exiting $[0,\infty)$ and the classical dynamical system defined by the piecewise linear mapping $x\mapsto\frac{1}{a}x+\frac{1}{2}\pmod{1}$. Namely, the trajectory of this system started at $X_{n}$ deterministically recovers the values of the killed chain in reversed time! We use this property to construct a suitable Banach space, where the transition operator of the killed chain has the compactness properties that let us apply a conventional argument of Perron–Frobenius type. The difficulty in finding such space stems from discreteness of the innovations.
@@
V. Wachtel was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project-ID 317210226 - SFB 1283.
1. Introduction
@@ 1.1. Motivation
where the starting point $X_{0}$ is independent of $\{\xi_{n}\}_{n\geq 1}$. This Markov chain is usually called an autoregressive sequence of order $1$, or AR($1$) in short. We will denote the distribution of $\{X_{n}\}_{n\geq 0}$ by $\mathbb{P}_{\nu}$ when $\nu$ is the distribution of $X_{0}$, or simply by $\mathbb{P}_{x}$ when $X_{0}=x$ is a fixed constant.
@@
as $n\to\infty$ for any starting point $x$. The proof of this result is based on a rather simple subadditivity argument, which allows one to prove existence of $\lambda_{a}$ but gives no information about any further properties of this exponent.
@@
The assumption that all moments of $\xi_{1}^{+}$ are finite is optimal, see Proposition 19 in [16]. On the contrary, the assumption on absolute continuity of the distribution of innovations was prompted by the method of the proof, which relied on compactness properties of the transition operator $P$ of the Markov chain $\{X_{n}\}$ killed at exiting $[0,\infty)$. To clarify, this operator that acts on measurable functions on $[0,\infty)$ by $Pf(x)=\mathbb{E}_{x}f(X_{1})\mathbbm{1}\{\tau>1\}$, and the killed chain is the sequence $\{X_{n}\}_{0\leq n<\tau}$.
@@
Initially we attempted a different approach, based on the observation that for every $a\in(0,\frac{2}{3})$, $\tau$ is the hitting time of zero for a non-negative Markov chain $\{Y_{n}\}$ obtained from $\{X_{n}\}$ by a certain aggregation of states. This aggregated (or lumped) chain has a finite number of states for almost every $a$. For such $a$, $\lambda_{a}$ is the Perron–Frobenius eigenvalue of the substochastic matrix corresponding to the transition operator of $\{Y_{n}\}$ killed at reaching 00. This approach is worth the attention even though it does not cover the remaining values of $a$ on the set of measure zero. We give the details in Section 2.3.
Unfortunately, neither of the two approaches works for arbitrary $a>\frac{2}{3}$ aside from some exceptions (see Remark 3). The case $a\in(0,\frac{1}{2}]$ can be solved by a simple direct argument, which gives that $\lambda_{a}=p$ for such $a$. Therefore, our main interest in $a\in(\frac{1}{2},\frac{2}{3}]$.
Let us comment on the background and the related literature besides [16]. The standard Perron–Frobenius method allows one to find the asymptotics of the probability that a finite state Markov chain avoids a subset of its states for a long time. Extending this to an infinite state space requires compactness properties of the transition operator of the corresponding killed chain. For persistence of general Markov chains, this is explored in [2], which gives further references and considers many examples including autoregressive chains with absolutely continuous innovations. A different approach is used in [1], which gives explicit combinatorial formulas for persistence probabilities of the AR($1$) chain with symmetric uniformly distributed innovations. The case where the innovations have a logarithmic tail is considered in [12]. For a general background on persistence problems, which have been extensively studied for many types of stochastic processes, we refer to the surveys [3, 9], where the second one gives a theoretical physics perspective.
@@ 1.2. Main results
It is easy to see that if the starting point $x$ is not greater than $1/(1-a)$, then all values of the chain $\{X_{n}\}$ do not exceed $1/(1-a)$. If the starting point $x$ is greater than $1/(1-a)$, then the trajectory of the chain is monotonically decreasing before the downcrossing of the level $1/(1-a)$. For this reason, it is natural to restrict our consideration to the starting points in the interval $[0,\frac{1}{1-a}]$, which we will always regard as the state space of the chain $X_{n}$ killed at the exit time $\tau$. (At the end of this section we shall mention how to generalise our main result to the case of arbitrary starting points.)
@@
We underline that $T_{a}$ is defined on the set $[0,\frac{1}{1-a}]\setminus I_{a}$, where
@@
We can also consider the case $a=\frac{2}{3}$, where we define $T_{2/3}$ as above but specify that $T_{2/3}(1)=0$ because (5) gives two values at $x=1$. Note that (6) ceases to hold for $a=\frac{2}{3}$: if $X_{n}=1$ and $\tau>n$, then we have two possible values for $X_{n-1}$, namely 00 and $3$.
Since $T_{a}$ is defined on $[0,\frac{1}{1-a}]\setminus I_{a}$, the iterates $T_{a}^{k}(x)$ are defined only up to the first hitting time of $I_{a}$, given by
$\varkappa_{a}(x)=\inf\left\{k\geq 0:\,T_{a}^{k}(x)\in I_{a}\right\}.$ (7)
@@
$\delta_{k}(x)=\mathbbm{1}\{T_{a}^{k}(x)<1\},\quad 0\leq k<\varkappa_{a}(x)+1,$
@@
$L_{0}(x)=0\quad\text{and}\quad L_{k}(x)=\sum_{i=0}^{k-1}\delta_{i}(x),\quad 1\leq k<\varkappa_{a}(x)+1.$
@@
To interpret the quantities $\delta_{k}(x)$ and $L_{k}(x)$, we note that the mapping $T_{a}$ is closely related to the mapping $\hat{T}_{a}$ given by $\hat{T}_{a}(x)=\frac{1}{a}x+\frac{1}{2}\pmod{1}$ for $0\leq x\leq 1$. Namely, since
$\hat{T}_{a}\Big{(}\frac{ax}{2}\Big{)}=\begin{cases}\frac{1}{2}(x+1),&0\leq x<1,\\ \frac{1}{2}(x-1),&1\leq x<3,\\ \frac{1}{2}(x-3),&3\leq x\leq\frac{2}{a},\\ \end{cases}$
we have $T_{a}(x)=\frac{2}{a}\hat{T}_{a}(\frac{ax}{2})$ for all $x$ in the domain of $T_{a}$ excluding $x=3$ when $a=\frac{2}{3}$. Consequently, for every $k\geq 1$ and $x\neq 3$ such that $\varkappa_{a}(x)\geq k$,
$T_{a}^{k}(x)=\frac{2}{a}\hat{T}_{a}^{k}\Big{(}\frac{ax}{2}\Big{)}.$ (8)
Note in passing that the sequences $\{\hat{T}_{a}^{k}(\frac{ax}{2})\}_{k\geq 0}$ with $x\in[0,\frac{1}{1-a}]$ can enter the set $[3,\frac{2}{a}]$ only through the “hole” $I_{a}$. The related dynamical systems defined by the $\beta$-transformations $x\mapsto\beta x\pmod{1}$ with a hole were studied in [11, 14].
On the other hand, $\hat{T}_{a}$ is one of the linear mod $1$ mappings $\hat{T}_{\beta,\alpha}(x)=\beta x+\alpha\pmod{1}$, where $\beta>1$ and $\alpha\in[0,1)$. By Parry [19, p. 101], every $y\in[0,1)$ can be written as
@@
where $d_{k}(y)$ are the “digits” of $y$ given by $d_{k}(y)=[\beta\hat{T}_{\beta,\alpha}^{k}(y)+\alpha]$, with []delimited-[][$\cdot$][ ⋅ ] denoting the integer part. This is a particular representation of $y$ in the base $\beta$, which we call the $(\beta,\alpha)$-expansion of $y$. Such expansions were first studied by Parry [19]. In the particularly important case $\alpha=0$ (and non-integer $\beta$), these are the so-called $\beta$-expansions introduced by Rényi [20].
[…]
The orbits of 00 and $1$ have a distinguished role for the linear mod $1$ transformations; for example, they appear in formula (58) for the invariant density of $\hat{T}_{\beta,\alpha}$. Similarly, the orbit of zero under $T_{a}$, that is $\{T_{a}^{k}(0):0\leq k<\varkappa_{a}(0)+1\}$, is very important for our analysis. For this reason, we denote
$\varkappa_{a}=\varkappa_{a}(0),\quad\delta_{k}=\delta_{k}(0),\quad L_{k}=L_{k}(0).$
@@
It is easy to see that if $\varkappa_{a}<\infty$, then the sequence $\{\hat{T}_{a}^{k}(0)\}_{k\geq\varkappa_{a}}$ strictly increases until it hits $[3a/2,1)$ at some moment $k$, hence $d_{k}(0)=2$. Then
@@
$x\in Q_{a}\text{ if and only if }\text{there are no $2$'s in the $\Big{(}\frac{1}{a},\frac{1}{2}\Big{)}$-expansion of }ax/2,$
once we define the digits of the $(\frac{3}{2},\frac{1}{2})$-expansion of $1$ as $0111\ldots$. Thus, $Q_{a}$ is fully analogous to the Cantor ternary set. Lastly, we note that in the case $a\in(\frac{2}{3},1)$, which is excluded from our consideration, the $(\frac{1}{a},\frac{1}{2})$-expansion of any point in $[0,1]$ has no $2$’s.
@@
Assume that $a\in(\frac{1}{2},\frac{2}{3}]$ and $p\in(0,1)$. Then there exists a constant $c\in(0,1)$ such that for every $x\in[0,\frac{1}{1-a}]$,
@@
$V(x)=\sum_{k=0}^{\varkappa_{a}}\Big{(}\frac{p}{\lambda_{a}}\Big{)}^{k}\Big{(}\frac{q}{p}\Big{)}^{L_{k}}\mathbbm{1}\{T_{a}^{k}(0)\leq x\}.$ (12)
The function $a\mapsto\lambda_{a}$ satisfies $\lambda_{1/2}=p$ and is continuous and non-decreasing on $[\frac{1}{2},\frac{2}{3}]$. More specifically, it is constant on every interval contained in $[\frac{1}{2},\frac{2}{3}]\setminus S$ and is constant on no open interval that intersects $S$, which is a closed set of Lebesgue measure zero. In other words, the Lebesgue–Stieltjes measure $d\lambda_{a}$ on $[\frac{1}{2},\frac{2}{3}]$ has no atoms, is singular, and its support is $S$.
Moreover, the conditional distributions converge weakly: for every $x,y\in[0,\frac{1}{1-a}]$,
@@
where the r.h.s. is the distribution function of a probability measure $\nu_{a}$ on $[1,1/(1-a)]$. This measure is quasi-stationary, i.e.
$\mathbb{P}_{\nu_{a}}(X_{1}\in A|\tau>1)=\nu_{a}(A),\quad A\in\mathcal{B}([0,1/(1-a)]).$
This measure has no atoms and is singular with respect to the Lebesgue measure except for the case $a=\frac{2}{3}$ and $p=\frac{1}{2}$, where $\nu_{2/3}$ is the uniform distribution on $[0,3]$. Lastly, the (topological) support of $\nu_{a}$ is the set $Q_{a}$ if there is no integer $k\geq 1$ such that $T^{k}_{a}(0)=\frac{1}{1-a}$, otherwise the support of $\nu_{a}$ is $Q_{a}\setminus\cup_{k=0}^{\infty}T_{a}^{-k}(0)$.
Let us make a few comments. We first present general explanations on the essence of this result, then give particular clarifying comments, and then list extensions and generalizations.
@@
    The key points of our proof are as follows. We will see that the killed transition operator $P$ is quasi-compact on a certain closed subspace of the Banach space $BV$ of functions of bounded variation on $[0,\frac{1}{1-a}]$, where $P$ can be represented by a linear operator acting on summable sequences of length $\varkappa_{a}^{\prime}+1$; see Proposition 12. We will use this representation to show that the leading eigenvalue of $P$ is $\lambda_{a}$ and the corresponding eigenfunction is $V$, and then prove (10). Our choice of the subspace is prompted by the mapping $T_{a}$, as explained in detail in Section 2.2. It turns out that $P$ is quasi-compact on the whole of $BV$; in particular, this implies weak convergence of the conditional distributions in (13).
    In all of our proofs it takes more effort to consider the case $p<1/2$, where we need to use uniform upper bounds for the frequencies $L_{n}(x)/n$ of zeros in the $(\frac{1}{a},\frac{1}{2})$-expansion of $ax/2$. We will obtain such bounds in Section 3.
@@
    The mapping $a\mapsto\lambda_{a}$ has intervals of constancy essentially due to discreteness of the innovations. In contrast, this mapping is strictly increasing if the innovations have a density that is strictly positive on $\mathbb{R}$ and log-concave, see [2, Theorem 2.7].
@@
    The fact that $\nu_{a}$ is singular continuous for every $a\in(1/2,2/3)$ and absolutely continuous for $a=2/3$ and $p=1/2$ reminds strongly of the same type of behaviour for Bernoulli convolutions. Recall that the Bernoulli convolution with parameter $a$ is the distribution $\pi_{a}$ of the series $\sum_{k=1}^{\infty}a^{k-1}\xi_{k}$. Note that $\pi_{a}$ is the stationary distribution for the chain $X_{n}$. It is well-known that if $p=1/2$, then $\pi_{a}$ is singular continuous for all $a<1/2$, $\pi_{1/2}$ is the uniform distribution on $[-2,2]$, and $\pi_{a}$ is absolutely continuous for almost all $a\in(1/2,1)$; see [23]. In Section 2.1 we shall describe the relation between our model and Bernoulli convolutions in more detail.

### v3

@@ Abstract.
We study the probability that an AR(1) Markov chain $X_{n+1}=aX_{n}+\xi_{n+1}$, where $a\in(0,1)$ is a constant, stays non-negative for a long time. We find the exact asymptotics of this probability and the weak limit of $X_{n}$ conditioned to stay non-negative, assuming that the i.i.d. innovations $\xi_{n}$ take only two values $\pm 1$ and $a\leq\frac{2}{3}$. This limiting distribution is quasi-stationary. It has no atoms and is singular with respect to the Lebesgue measure when $\frac{1}{2}<a\leq\frac{2}{3}$, except for the case $a=\frac{2}{3}$ and $\mathbb{P}(\xi_{n}=1)=\frac{1}{2}$, where this distribution is uniform on the interval $[0,3]$. This is similar to the properties of Bernoulli convolutions. For $0<a\leq\frac{1}{2}$, the situation is much simpler, and the limiting distribution is a $\delta$-measure.
To prove these results, we uncover a close connection between $X_{n}$ killed at exiting $0,\infty)$ and the classical dynamical system defined by the piecewise linear mapping $x\mapsto\frac{1}{a}x+\frac{1}{2}\pmod{1}$. Namely, the trajectory of this system started at $X_{n}$ deterministically recovers the values of the killed chain in reversed time. We use this fact to construct a suitable Banach space, where the transition operator of the killed chain has the compactness properties that allow us to apply a conventional argument of the Perron–Frobenius type.
@@
V. Vysotsky was supported in part by Dr Perry James (Jim) Browne Research Centre. V. Wachtel was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project-ID 317210226 - SFB 1283.
1. Introduction and main results
@@ 1.1. Motivation
where the starting point $X_{0}$ is independent of $\{\xi_{n}\}_{n\geq 1}$. This Markov chain is usually called an autoregressive sequence of order $1$, or AR($1$) in short. We will denote the distribution of $\{X_{n}\}_{n\geq 0}$ by $\mathbb{P}_{\nu}$ when $\nu$ is the distribution of $X_{0}$, or simply by $\mathbb{P}_{x}$ when $X_{0}=x$ is a fixed real number.
@@
as $n\to\infty$ for any starting point $x>0$. The proof of this result is based on a rather simple subadditivity argument, which allows one to prove existence of $\lambda_{a}$ but gives no information about any further properties of this exponent.
@@
The assumption that all moments of $\xi_{1}^{+}$ are finite is optimal, see Proposition 19 in [17]. On the contrary, the assumption on absolute continuity of the distribution of innovations was prompted by the method of the proof, which relied on compactness properties of the transition operator $P$ of the Markov chain $\{X_{n}\}$ killed at exiting $0,\infty)$. To clarify, this operator that acts on measurable functions on $0,\infty)$ by $Pf(x)=\mathbb{E}_{x}f(X_{1})\mathbbm{1}\{\tau>1\}$, and the killed chain is the sequence $\{X_{n}\}_{0\leq n<\tau}$.
@@
Initially we attempted a different approach, based on the observation that for every $a\in(0,\frac{2}{3})$, $\tau$ is the hitting time of zero for a non-negative Markov chain $\{Y_{n}\}$ obtained from $\{X_{n}\}$ by a certain aggregation of states. This aggregated (or lumped) chain has a finite number of states for almost every $a$. For such $a$, $\lambda_{a}$ is the Perron–Frobenius eigenvalue of the substochastic matrix corresponding to the transition operator of $\{Y_{n}\}$ killed at reaching 0. This approach is worth the attention even though it does not cover the remaining values of $a$ on the set of measure zero. We give the details in Section 2.2.
Unfortunately, neither of the two approaches works for arbitrary $a>\frac{2}{3}$ aside some exceptional values (see Remark 9). The case $a\in(0,\frac{1}{2}]$ can be solved by a simple direct argument, which gives that $\lambda_{a}=p$ for such $a$. Therefore, our main interest in $a\in(\frac{1}{2},\frac{2}{3}]$.
Let us comment on the background and the related literature besides [17]. The standard Perron–Frobenius method allows one to find the asymptotics of the probability that a finite state Markov chain avoids a subset of its states for a long time. Extending this to an infinite state space requires compactness properties of the transition operator of the corresponding killed chain. For persistence of general Markov chains, this is explored in [3], which gives further references and considers many examples including autoregressive chains with absolutely continuous innovations. A different approach is used in [1], which gives explicit combinatorial formulas for persistence probabilities of the AR($1$) chain with symmetric uniformly distributed innovations. The case where the innovations have logarithmic tail probabilities is considered in [13]. For a general background on persistence problems, which have been extensively studied for many types of stochastic processes, we refer to the surveys [4, 10], where the second one gives a theoretical physics perspective.
@@ 1.2. Main results
It is easy to see that if the starting point $x$ is not greater than 1/(1a)1/(1-a), then all values of the chain $\{X_{n}\}$ do not exceed 1/(1a)1/(1-a). If the starting point $x$ is greater than 1/(1a)1/(1-a), then the trajectory of the chain is monotonically decreasing before the downcrossing of the level 1/(1a)1/(1-a). For this reason, it is natural to restrict our consideration to the starting points in the interval $0,\frac{1}{1-a}]$, which we will always regard as the state space of the chain $\{X_{n}\}$ killed at the exit time $\tau$. We stress that our results can be easily generalized to arbitrary starting points $x\geq 0$; see, for example, Corollary 2 below.
@@
We underline that $T_{a}$ is defined on the set $0,\frac{1}{1-a}]\setminus I_{a}$, where
@@
We can also consider the case $a=\frac{2}{3}$, where we define $T_{2/3}$ as above but specify that $T_{2/3}(1)=0$ because (5) gives two values at $x=1$. Note that (6) ceases to hold for $a=\frac{2}{3}$: if $X_{n}=1$ and $\tau>n$, then we have two possible values for $X_{n-1}$, namely 0 and $3$.
Since $T_{a}$ is defined on $0,\frac{1}{1-a}]\setminus I_{a}$, the iterates $T_{a}^{k}(x)$ are defined only up to the first hitting time of $I_{a}$, given by
$\varkappa_{a}(x)=\inf\left\{k\geq 0:\,T_{a}^{k}(x)\in I_{a}\right\}\in[0,\infty].$ (7)
@@
$\delta_{k}(x)=\mathbbm{1}\{T_{a}^{k}(x)<1\},\quad 0\leq k<\varkappa_{a}(x)+1,$ (8)
@@
$L_{0}(x)=0\quad\text{and}\quad L_{k}(x)=\sum_{i=0}^{k-1}\delta_{i}(x),\quad 1\leq k<\varkappa_{a}(x)+1.$ (9)
@@
To interpret the quantities $\delta_{k}(x)$ and $L_{k}(x)$, we note that the mapping $T_{a}$ is closely related to the mapping $\widehat{T}_{a}$ given by $\widehat{T}_{a}(x)=\frac{1}{a}x+\frac{1}{2}\pmod{1}$ for $0\leq x\leq 1$. Namely, since
[…]
we have $T_{a}(x)=\frac{2}{a}\widehat{T}_{a}(\frac{ax}{2})$ for all $x$ in the domain of $T_{a}$ excluding $x=3$ when $a=\frac{2}{3}$. Consequently, for every $k\geq 1$ and $x\neq 3$ such that $\varkappa_{a}(x)\geq k$,
$T_{a}^{k}(x)=\frac{2}{a}\widehat{T}_{a}^{k}\Big(\frac{ax}{2}\Big).$ (10)
Note in passing that the sequences $\{\frac{2}{a}\widehat{T}_{a}^{k}(\frac{ax}{2})\}_{k\geq 0}$ with $x\in[0,\frac{1}{1-a}]$ can enter the set $3,\frac{2}{a}]$ only through the “hole” $I_{a}$. The related dynamical systems defined by the $\beta$-transformations $x\mapsto\beta x\pmod{1}$ with a hole were studied in [12, 15].
On the other hand, $\widehat{T}_{a}$ is one of the linear mod $1$ mappings $\widehat{T}_{\beta,\alpha}(x)=\beta x+\alpha\pmod{1}$, where $\beta>1$ and $\alpha\in[0,1)$. By Parry [21, p. 101], every $y\in[0,1)$ can be written as
@@
where $d_{k}(y)$ are the “digits” of $y$ given by $d_{k}(y)=[\beta\widehat{T}_{\beta,\alpha}^{k}(y)+\alpha]$, with $\cdot$] denoting the integer part. This is a particular representation of $y$ in the base $\beta$, which we call the $(\beta,\alpha)$-expansion of $y$. Such expansions were first studied by Parry [21]. In the particularly important case $\alpha=0$ (and non-integer $\beta$), these are the so-called $\beta$-expansions introduced by Rényi [22].
It follows from (10) that $1-\delta_{k}(x)$, for $0\leq k<\varkappa_{a}(x)+1$, are the first digits in the $(\frac{1}{a},\frac{1}{2})$-expansion of $ax/2$ (unless $x=3$), and thus $L_{k}(x)$ is the number of 0’s in the first $k$ digits of this expansion. Moreover, we will also show that $\delta_{0}(x),\delta_{1}(x),\ldots$ for $a=2/3$ are the digits of the $\frac{3}{2}$-expansion of 1x/31-x/3 (up to a minor modification); see (78) below, where we write $T_{2/3}$ in terms of the $\frac{3}{2}$-transformation $\widehat{T}_{3/2,\,0}$. There are many works on digit frequencies in $\beta$-expansions, including [8, 24]; unfortunately, they hardly consider concrete values of $(\beta,x)$.
The orbits of 0 and $1$ have a distinguished role for the linear mod $1$ transformations; for example, they appear in formula (68) for the invariant density of $\widehat{T}_{\beta,\alpha}$. Similarly, the orbit of zero under $T_{a}$, that is $\{T_{a}^{k}(0):0\leq k<\varkappa_{a}(0)+1\}$, is very important for our analysis. For this reason, we denote
$\varkappa_{a}=\varkappa_{a}(0),\quad\delta_{k}=\delta_{k}(0),\quad L_{k}=L_{k}(0).$ (12)
@@
It is easy to see that if $\varkappa_{a}<\infty$, then the sequence $\{\widehat{T}_{a}^{k}(0)\}_{k\geq\varkappa_{a}}$ strictly increases until it hits $[3a/2,1)$ at some moment $k^{\prime}$, hence $d_{k^{\prime}}(0)=2$. Then
@@
[…]
once we re-define the digits of the $(\frac{3}{2},\frac{1}{2})$-expansion of $1$ as $0111\ldots$. Thus, $Q_{a}$ is fully analogous to the Cantor ternary set. Lastly, we note that in the case $a\in(\frac{2}{3},1)$, which is excluded from our consideration, the $(\frac{1}{a},\frac{1}{2})$-expansion of any point in $[0,1]$ has no $2$’s.
@@
Let $\{X_{n}\}$ be a Markov chain defined by equation (1) with some $a\in(\frac{1}{2},\frac{2}{3}]$. Assume that the innovations $\{\xi_{n}\}$ satisfy (4) with some $p\in(0,1)$. Then there exists a constant $c\in(0,1)$ such that, uniformly in $x\in[0,\frac{1}{1-a}]$, we have
@@
$V(x)=\sum_{k=0}^{\varkappa_{a}}\Big(\frac{p}{\lambda_{a}}\Big)^{k}\Big(\frac{q}{p}\Big)^{L_{k}}\mathbbm{1}\{T_{a}^{k}(0)\leq x\},$ (16)
with $\delta_{k}$ and $L_{k}$ defined in (8), (9), and (12). The constant $c$ is given in (40) below.
The function $a\mapsto\lambda_{a}$ satisfies $\lambda_{1/2}=p$ and is continuous and non-decreasing on $\frac{1}{2},\frac{2}{3}]$. More specifically, it is constant on every interval contained in $\frac{1}{2},\frac{2}{3}]\setminus S$ and is constant on no open interval intersecting $S$, which is a closed set of Lebesgue measure zero defined in (13). In other words, the Lebesgue–Stieltjes measure $d\lambda_{a}$ on $\frac{1}{2},\frac{2}{3}]$ has no atoms, is singular, and its topological support is $S$.
Moreover, the conditional distributions converge weakly, uniformly in $x,y\in[0,\frac{1}{1-a}]$:
@@
where the right-hand side is the distribution function of a probability measure $\nu_{a}$ on $0,\frac{1}{1-a}]$. This measure is quasi-stationary, i.e.
$\mathbb{P}_{\nu_{a}}(X_{1}\in A|\tau>1)=\nu_{a}(A),\quad A\in\mathcal{B}([0,1/(1-a)]).$ (18)
This measure has no atoms and is singular with respect to the Lebesgue measure, except in the case $a=\frac{2}{3}$ and $p=\frac{1}{2}$, where $\nu_{2/3}$ is the uniform distribution on $[0,3]$. The topological support of $\nu_{a}$ is the set of non-isolated points of $Q_{a}$, that is, $Q_{a}$ itself if there is no integer $k\geq 1$ such that $T^{k}_{a}(0)=\frac{1}{1-a}$, and otherwise $Q_{a}\setminus\cup_{k=0}^{\infty}T_{a}^{-k}(0)$.
Corollary 2.
Under the assumptions of Theorem 1, for every $x\geq 0$, we have
$\mathbb{P}_{x}(\tau>n)\sim c\lambda_{a}^{n}\mathbb{E}_{x}[\lambda_{a}^{-\sigma}V(X_{\sigma})]$
where
$\sigma:=\inf\left\{n\geq 0:X_{n}\leq 1/(1-a)\right\}.$
Moreover, the weak convergence (17) holds true for all $x\geq 0$.
Fig. 1. The graphs of $\lambda_{a}(p)$ for fixed $p$.
Fig. 2. The graph of $\varkappa_{a}^{\prime}$.
a) $a=0.55$
b) $a=0.6$
c) $a=0.65$
d) $a=2/3$
Fig. 3. The distribution functions of $\nu_{a}$ for $p=0.3$, $p=0.5$, $p=0.7$ (top to bottom).
Let us make a few comments.
@@
    The value of $\lambda_{a}$ is rather explicit from (15) when $\varkappa_{a}^{\prime}<\infty$; see Figure 1.2. In this case the left-hand side in (15) can be written as a finite sum even if $\varkappa_{a}=\infty$ because then the sequence $\{\delta_{k}\}$ is eventually periodic. Thus, $\lambda_{a}$ is a solution to a polynomial equation of order $\varkappa_{a}^{\prime}+1$, and we will give some of its values in Section 2.2. See Figure 1.2 for a graph of $\varkappa_{a}^{\prime}$. If $\varkappa_{a}^{\prime}=\infty$, no simplification of (15) appears to be possible because of chaotic behaviour of the orbit of zero, unless $a=\frac{2}{3}$ and $p=\frac{1}{2}$. In this case
    $\lambda_{2/3}(1/2)=3/4;$
    see Proposition 10. We found this value computing the left-hand side of (15) for $\lambda=ap$ using the $(\frac{1}{a},\frac{1}{2})$-expansion of 0. In Section 2.4, we will present an alternative way to compute $\lambda_{2/3}(1/2)$ and to establish (14) with explicit expressions for $c$ and $V(x)$ in the case $a=\frac{2}{3}$ and $p=\frac{1}{2}$, based on a close connection between the transition operator of $\{X_{n}\}$ killed at leaving $0,\infty)$ and the transfer operator associated with $T_{2/3}$.
@@
    The mapping $a\mapsto\lambda_{a}$ has intervals of constancy essentially due to discreteness of the innovations. These intervals are the intervals of constancy of $a\mapsto\varkappa_{a}^{\prime}$, cf.  Figures 1.2 and 1.2. In contrast, the mapping $a\mapsto\lambda_{a}$ is strictly increasing if the innovations have a density that is strictly positive on $\mathbb{R}$ and log-concave, see Aurzada et al. [3, Theorem 2.7]. In the particular case of the standard normal innovations, Aurzada and Kettner [2] derived a series expansion for $\lambda_{a}$. For the uniformly distributed innovations, $\lambda_{a}$ was found in a rather explicit form by Alsmeyer et al. [1], see Propositions 2.4 and 3.11, and Remark 5.9(b) there. Based on numerical evidence, it appears that $a\mapsto\lambda_{a}$ is strictly increasing in this case as well.
@@
    The fact that $\nu_{a}$ is singular continuous for every $a\in(\frac{1}{2},\frac{2}{3})$ and absolutely continuous for $a=\frac{2}{3}$ and $p=\frac{1}{2}$ reminds us strongly of the same type of behaviour for Bernoulli convolutions. Recall that the Bernoulli convolution with parameter $a$ is the distribution $\pi_{a}$ of the series $\sum_{k=1}^{\infty}a^{k-1}\xi_{k}$. This is the stationary distribution for the chain $X_{n}$. It is well-known that if $p=\frac{1}{2}$, then $\pi_{a}$ is singular continuous for all $a<\frac{1}{2}$, $\pi_{1/2}$ is the uniform distribution on [2,2][-2,2], and $\pi_{a}$ is absolutely continuous for almost all $a\in(\frac{1}{2},1)$; see [25]. In Section 2.1 we shall describe the relation between our model and Bernoulli convolutions in more detail.